# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 10 Isomorphism of $X_{p}$ with $Y_{p}$
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. In the latter case $G$ means $\mathbb{G}$.
Theorem 11.5.
For $v\in W$, denote $\sf X_{v}$ to be the strictly full subcategory of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ generated by $\{M\otimes X_{v}\mid M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\}$. Then the category $\mathcal{D}$ has a $G$-linear semiorthogonal decomposition
$\mathcal{D}=\langle\sf X_{v}\rangle_{v\in W}$ (44)
with respect to the order $\prec$ on the Weyl group $W$. Each subcategory $\sf X_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
Proof.
Denote $\Phi_{p}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathcal{D}$ the $G$-linear functor $M\rightarrow M\otimes_{{\sf k}}X_{p}$. Now Corollary 11.4 and Proposition 8.8 give that $\Phi_{p}$ is fully faithful.
Let now $w\succ v$. By Corollary 11.4 and [29, I Proposition 3.6] we have ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{w}\otimes M)={\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{w})\otimes M=0$ for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. By Proposition 11.6 below, each subcategory ${\sf X}_{v},v\in W$ is admissible in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Proposition 8.7 then gives that the sequence $\langle\sf X_{v}\rangle_{v\in W}$ of admissible subcategories of $\mathcal{D}$ is $G$-semiorthogonal with respect to the order $\prec$ on $W$. Lemma 9.2 states that the triangulated hull of $\langle\sf X_{v}\rangle_{v\in W}$ coincides with $\mathcal{D}$. Thus, it is a $G$-linear semi-orthogonal decomposition of $\mathcal{D}$. ∎
@@
For each $p\in W$ the subcategory $\sf X_{v}$ is an admissible subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$.
@@
$-\;\otimes_{\sf k}-:{\mathrm{rep}}_{\mathrm{fr}}(B)\times{\mathrm{rep}}_{\mathrm{fr}}(B)\to{\mathrm{rep}}_{\mathrm{fr}}(B)$
@@
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):{\mathrm{rep}}_{\mathrm{fr}}(B)\times{\mathrm{rep}}_{\mathrm{fr}}(B)\to{\mathrm{rep}}_{\mathrm{fr}}(B)$
to the bounded derived category of ${\mathrm{rep}}(B)$. We put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$. Then the right adjoint of the inclusion of $\sf X_{v}$ is
${\mathrm{Rind}}_{B}^{G}(-\otimes_{\sf k}X_{p}^{*})\otimes_{\sf k}X_{p}$
@@
$({\mathrm{Rind}}_{B}^{G}((-)^{*}\otimes_{\sf k}X_{p}))^{*}\otimes_{\sf k}X_{p}.$
12 Full exceptional collections in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/B)$
We recall first the sheafification functor $\mathcal{L}$. Let $G$ be a flat $\sf k$-group scheme acting freely (from the right) on a flat $\sf k$-scheme $X$ such that $X/G$ is a scheme. Associated to each $G$-module $M$ is a sheaf $\mathcal{L}(M)=\mathcal{L}_{X/G}(M)$ on $X/G$, [29, I, Section 5.8]. The functor $\mathcal{L}:{{\mathrm{Rep}}}({G})\rightarrow{\rm Sh}(X/G)$ is exact and lands in the subcategory ${\mathsf{QCoh}}(X/G)$ of quasi–coherent $\mathcal{O}_{X/G}$-modules of ${\rm Sh}(X/G)$, [29, I, Proposition 5.9]. Since $\mathcal{L}$ is exact, denote also $\mathcal{L}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits({{\mathrm{Rep}}}({G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits_{{\mathsf{QCoh}}(X/G)}{\rm Sh}(X/G)$ [^3] The derived category of complexes on ${\rm Sh}(X/G)$ with quasi-coherent cohomology. its derived functor.
Now set $G:=B,X=G$ (so $X=G/B$) in the definition of the functor $\mathcal{L}$. It restricts to the functor between the bounded derived categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}({B}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/B)$. We put ${\mathcal{X}}_{v}:=\mathcal{L}(X_{v})$ for $v\in W$.
[…]
@@
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/B)$. Let $v,w\in W$.
@@
$(1)$. By Proposition 8.6, the functor $\mathop{\mathrm{RHom}}\nolimits_{G/B}(\mathcal{L}(-),\mathcal{L}(-))$ canonically factors through ${\mathrm{Rind}}_{B}^{G}((-)^{\ast}\otimes(-))$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}({\mathcal{X}}_{v},{\mathcal{X}}_{v})=\mathop{\mathrm{RHom}}\nolimits_{G/B}(\mathcal{L}(X_{v}),\mathcal{L}(X_{v}))={\sf For}\circ{\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})$ where $\sf For$ is the forgetful functor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\sf k}\textrm{-}\mathrm{mod}})$. This also holds when ${\sf k}={\mathbb{Z}}$. Now by Corollary 11.4 (1), ${\mathrm{Rind}}_{B}^{G}(X_{v}^{\ast}\otimes X_{v})={\sf k}$.
@@
By Lemma 9.2, ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$ coincides with the hull of ${\mathrm{rep}}(B)$ in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. In particular, the abelian category ${\mathrm{rep}}(B)\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ is contained in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$. By definition, ${\mathcal{X}}_{v}=\mathcal{L}(X_{v})$; applying the functor $\mathcal{L}$ to the above inclusion, we see that the category of $G$-equivariant coherent sheaves ${\sf Coh}^{G}(G/B)=\mathcal{L}({\mathrm{rep}}(B))$ is contained in the image under $\mathcal{L}$ of ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$. Recall that the functor $\mathcal{L}$ is monoidal; thus, $\mathcal{L}(\nabla_{\lambda}\otimes X_{v})=\mathcal{L}(\nabla_{\lambda})\otimes{\mathcal{X}}_{v}$. Now $\mathcal{L}(\nabla_{\lambda})$ is a ($G$-equivariant) trivial vector bundle on $G/B$, so the image of ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$ under $\mathcal{L}$ coincides with the hull $\langle{\mathcal{X}}_{v},v\in W\rangle$ of $\{{\mathcal{X}}_{v},v\in W\}$. We conclude that $\langle{\mathcal{X}}_{v},v\in W\rangle$ contains the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{G}(G/B)$.
Now since $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{G}(G/B)\subset\langle{\mathcal{X}}_{v},v\in W\rangle$, the category $\langle{\mathcal{X}}_{v},v\in W\rangle$ contains all line bundles $\mathcal{L}_{\lambda},\lambda\in X(T)_{+}$, the group $G$ being simply connected. Now $\langle{\mathcal{X}}_{v},v\in W\rangle^{\perp}=0$ by Corollary 8.11. By Lemma 7.1 we get $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))=\langle{\mathcal{X}}_{v},v\in W\rangle$. ∎

### v3

@@ 10 Isomorphism of $X_{p}$ with $Y_{p}$
Let $p\in W$. The strictly full subcategory of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ generated by $\{M\otimes X_{p}\mid M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\}$ is an admissible subcategory.
@@
$-\;\otimes_{\sf k}-:{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})\times{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})\to{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})$
@@
$\mathop{\mathrm{Hom}}\nolimits_{\sf k}(-,-):{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})\times{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})\to{\mathrm{rep}}_{\mathrm{fr}}(\mathbf{B})$
to the bounded derived category of ${\mathrm{rep}}(\mathbf{B})$. We put $M^{*}=\mathop{\mathrm{Hom}}\nolimits_{\sf k}(M,{\sf k})$. Then the right adjoint of the inclusion of the subcategory is
${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(-\otimes_{\sf k}X_{p}^{*})\otimes_{\sf k}X_{p}$
@@
$({\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}((-)^{*}\otimes_{\sf k}X_{p}))^{*}\otimes_{\sf k}X_{p}.$
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. In the latter case $\mathbf{G}$ means $\mathbb{G}$.
@@
For $v\in W$, denote $\sf X_{v}$ to be the strictly full subcategory of $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ generated by $\{M\otimes X_{v}\mid M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\}$. Then the category $\mathcal{D}$ has a $\mathbf{G}$-linear semiorthogonal decomposition
$\mathcal{D}=\langle\sf X_{v}\rangle_{v\in W}$ (47)
with respect to the order $\prec$ on the Weyl group $W$. Each subcategory $\sf X_{v}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$.
Proof.
Denote $\Phi_{p}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\rightarrow\mathcal{D}$ the $\mathbf{G}$-linear functor $M\rightarrow M\otimes_{{\sf k}}X_{p}$. Now Corollary 11.4 and Proposition 8.11 give that $\Phi_{p}$ is fully faithful.
Let now $w\succ v$. By Corollary 11.4 and [32, I Proposition 3.6] we have ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(X_{v}^{\ast}\otimes X_{w}\otimes M)={\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(X_{v}^{\ast}\otimes X_{w})\otimes M=0$ for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$. By Proposition 11.5, each subcategory ${\sf X}_{v},v\in W$ is admissible in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. Proposition 8.9 then gives that the sequence $\langle\sf X_{v}\rangle_{v\in W}$ of admissible subcategories of $\mathcal{D}$ is $\mathbf{G}$-semiorthogonal with respect to the order $\prec$ on $W$. Lemma 9.4 states that the triangulated hull of $\langle\sf X_{v}\rangle_{v\in W}$ coincides with $\mathcal{D}$. Thus, it is a $\mathbf{G}$-linear semi-orthogonal decomposition of $\mathcal{D}$. ∎
12 Full exceptional collections in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}/\mathbf{B})$
We recall first the sheafification functor $\mathcal{L}$. Let $\mathbf{G}$ be a flat $\sf k$-group scheme acting freely (from the right) on a flat $\sf k$-scheme $X$ such that $X/\mathbf{G}$ is a scheme. Associated to each $\mathbf{G}$-module $M$ is a sheaf $\mathcal{L}(M)=\mathcal{L}_{X/\mathbf{G}}(M)$ on $X/\mathbf{G}$, [32, I, Section 5.8]. The functor $\mathcal{L}:{{\mathrm{Rep}}}({\mathbf{G}})\rightarrow{\rm Sh}(X/\mathbf{G})$ is exact and lands in the subcategory ${\mathsf{QCoh}}(X/\mathbf{G})$ of quasi–coherent $\mathcal{O}_{X/\mathbf{G}}$-modules of ${\rm Sh}(X/\mathbf{G})$, [32, I, Proposition 5.9]. Since $\mathcal{L}$ is exact, denote also $\mathcal{L}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits({{\mathrm{Rep}}}({\mathbf{G}}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits_{{\mathsf{QCoh}}(X/\mathbf{G})}{\rm Sh}(X/\mathbf{G})$ [^2] The derived category of complexes on ${\rm Sh}(X/\mathbf{G})$ with quasi-coherent cohomology. its derived functor.
Now set $\mathbf{G}:=\mathbf{B},X=\mathbf{G}$ (so $X=\mathbf{G}/\mathbf{B}$) in the definition of the functor $\mathcal{L}$. It restricts to the functor between the bounded derived categories $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\mathrm{rep}}}({\mathbf{B}}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}/\mathbf{B})$. We put ${\mathcal{X}}_{v}:=\mathcal{L}(X_{v})$ for $v\in W$.
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. In the latter case $\mathbf{G}$ means $\mathbb{G}$. Recall that $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}/\mathbf{B}):=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$. Theorem 11.6 implies the following:
Theorem 12.1.
Let $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}/\mathbf{B})$. Let $v,w\in W$.
@@
$(1)$. By Proposition 8.6, the functor $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}/\mathbf{B}}(\mathcal{L}(-),\mathcal{L}(-))$ canonically factors through ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}((-)^{\ast}\otimes(-))$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathcal{D}}({\mathcal{X}}_{v},{\mathcal{X}}_{v})=\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}/\mathbf{B}}(\mathcal{L}(X_{v}),\mathcal{L}(X_{v}))={\sf For}\circ{\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(X_{v}^{\ast}\otimes X_{v})$ where $\sf For$ is the forgetful functor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({{\sf k}\textrm{-}\mathrm{mod}})$. This also holds when ${\sf k}={\mathbb{Z}}$. Now by Corollary 11.4 (1), ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}(X_{v}^{\ast}\otimes X_{v})={\sf k}$.
@@
$(3).$ We will prove that the triangulated hull $\langle{\mathcal{X}}_{v},v\in W\rangle$ of $\{{\mathcal{X}}_{v},v\in W\}$ in $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$ is the whole $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$. Observe that by the previous items (1) and (2), the collection of ${\mathcal{X}}_{v},v\in W$ is exceptional. Thus, $\langle{\mathcal{X}}_{v},v\in W\rangle\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$ is an admissible (hence thick) subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$, see Section 7. We are bound to show that $\langle{\mathcal{X}}_{v},v\in W\rangle^{\perp}=0$. Let us show that $\langle{\mathcal{X}}_{v},v\in W\rangle$ contains the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{\mathbf{G}}(\mathbf{G}/\mathbf{B}):=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\sf Coh}^{\mathbf{G}}(\mathbf{G}/\mathbf{B}))$ of all $\mathbf{G}$-equivariant objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathop{\sf{Coh}}\nolimits(\mathbf{G}/\mathbf{B}))$.
By Lemma 9.4, ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$ coincides with the hull of ${\mathrm{rep}}(\mathbf{B})$ in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$. In particular, the abelian category ${\mathrm{rep}}(\mathbf{B})\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$ is contained in ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$. By definition, ${\mathcal{X}}_{v}=\mathcal{L}(X_{v})$; applying the functor $\mathcal{L}$ to the above inclusion, we see that the category of $\mathbf{G}$-equivariant coherent sheaves ${\sf Coh}^{\mathbf{G}}(\mathbf{G}/\mathbf{B})=\mathcal{L}({\mathrm{rep}}(\mathbf{B}))$ is contained in the image under $\mathcal{L}$ of ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$. Recall that the functor $\mathcal{L}$ is monoidal; thus, $\mathcal{L}(\nabla_{\lambda}\otimes X_{v})=\mathcal{L}(\nabla_{\lambda})\otimes{\mathcal{X}}_{v}$. Now $\mathcal{L}(\nabla_{\lambda})$ is a ($\mathbf{G}$-equivariant) trivial vector bundle on $\mathbf{G}/\mathbf{B}$, so the image of ${\mathrm{hull}}(\{\nabla_{\lambda}\otimes X_{v}\}_{v\in W})$ under $\mathcal{L}$ coincides with the hull $\langle{\mathcal{X}}_{v},v\in W\rangle$ of $\{{\mathcal{X}}_{v},v\in W\}$. We conclude that $\langle{\mathcal{X}}_{v},v\in W\rangle$ contains the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{\mathbf{G}}(\mathbf{G}/\mathbf{B})$.
Now since $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{\mathbf{G}}(\mathbf{G}/\mathbf{B})\subset\langle{\mathcal{X}}_{v},v\in W\rangle$, the category $\langle{\mathcal{X}}_{v},v\in W\rangle$ contains all line bundles $\mathcal{L}_{\lambda},\lambda\in X(\mathbf{T})_{+}$, the group $\mathbf{G}$ being simply connected. Now $\langle{\mathcal{X}}_{v},v\in W\rangle^{\perp}=0$ by Corollary 8.15. By Lemma 7.5 we get $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))=\langle{\mathcal{X}}_{v},v\in W\rangle$. ∎
Remark 12.2.
Theorem 12.1 is in fact a simple instance of base change for semiorthogonal decompositions, [36], which in this particular case can be stated in elementary terms. In more detail, consider a flat morphism $\pi\colon X\rightarrow S$ between smooth projective varieties, a closed point $s\in S$ and the base change diagram along the embedding $i_{s}\colon s\hookrightarrow S$:
@@
Assume given an $S$-linear semiortogonal decomposition $\langle\mathcal{D}_{1},\dots,\mathcal{D}_{n}\rangle$ of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$, such that each admissible subcategory $\mathcal{D}_{i}$ is equivalent to $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(S)$. Then each embedding functor $\Phi_{k}:\mathcal{D}_{k}\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X),k=1,\dots,n$ is given by $\Phi_{k}(-)=(-)\otimes E_{k}$ where $E_{k},k=1,\dots n$ is a collection of objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ with the following two properties: 1) $R\pi_{\ast}R{\mathcal{H}om}(E_{i},E_{j})=0$ for $i>j$ and 2) $R\pi_{\ast}{\mathcal{H}om}(E_{k},E_{k})=\mathcal{O}_{S}$ for all $k$ (cf. Propositions 8.9 and 8.11). Restricting the objects $E_{k}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ to the fiber $X_{s}$, one obtains a collection of objects $\mathcal{E}_{k}:=i^{\ast}E_{k},k=1,\dots n$. Now the claim is that the collection $\mathcal{E}_{k},k=1,\dots n$ is exceptional in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X_{s})$: for instance, to see that $\mathop{\mathrm{RHom}}\nolimits_{X_{s}}(\mathcal{E}_{i},\mathcal{E}_{j})=0$ for $i>j$, by [34, Lemma 2.32] base change holds for the above Cartesian square, thus
$\mathop{\mathrm{RHom}}\nolimits_{X_{s}}(\mathcal{E}_{i},\mathcal{E}_{j})=R{\pi_{s}}_{\ast}(\mathcal{E}_{i}^{\ast}\otimes^{\mathbb{L}}\mathcal{E}_{j})=R{\pi_{s}}_{\ast}i^{\ast}(E_{i}^{\ast}\otimes^{\mathbb{L}}E_{j})=$ (48)
$i^{\ast}_{s}R\pi_{\ast}(E_{i}^{\ast}\otimes^{\mathbb{L}}E_{j})=i^{\ast}_{s}R\pi_{\ast}R{\mathcal{H}om}(E_{i},E_{j})=0$ (49)
by 1) above. Similarly for $\mathop{\mathrm{RHom}}\nolimits_{X_{s}}(\mathcal{E}_{i},\mathcal{E}_{i})={\sf k}$ for $i=1,\dots n$. Finally, the objects $\mathcal{E}_{i},i=1,\dots n$ generate $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X_{s})$: for if there was a non-trivial object $\mathcal{F}\in\langle\mathcal{E}_{1},\dots,\mathcal{E}_{n}\rangle^{\perp}$, its pushforward ${i_{s}}_{\ast}\mathcal{F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ would be a non-trivial object in the right orthogonal to the semiortogonal decomposition $\langle\mathcal{D}_{1},\dots,\mathcal{D}_{n}\rangle$ of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$, a contradiction.